OpenStudy (anonymous):

Convert 30 liters per hour to gallon per hour?

OpenStudy (adajiamcneal):

7.925 or ~8 gals

OpenStudy (anonymous):

7.92516157 US gallon per hour

OpenStudy (anonymous):

what she said

OpenStudy (anonymous):

Thanks guys!

OpenStudy (mathstudent55):

\(1 ~gal \approx 3.785~liter\) \(30 \dfrac{liter}{hour} \times \dfrac{gal}{3.785~liter} = 7.926 \dfrac{gal}{hour}\)